# Learning Multi-agent Multi-machine Tending by Mobile Robots

### v1

@@ Abstract
Robotics can help address the growing worker shortage challenge of the manufacturing industry. As such, machine tending is a task collaborative robots can tackle that can also highly boost productivity. Nevertheless, existing robotics systems deployed in that sector rely on a fixed single-arm setup, whereas mobile robots can provide more flexibility and scalability. In this work, we introduce a multi-agent multi-machine tending learning framework by mobile robots based on Multi-agent Reinforcement Learning (MARL) techniques with the design of a suitable observation and reward. Moreover, an attention-based encoding mechanism is developed and integrated into Multi-agent Proximal Policy Optimization (MAPPO) algorithm to boost its performance for machine tending scenarios. Our model (AB-MAPPO) outperformed MAPPO in this new challenging scenario in terms of task success, safety, and resources utilization. Furthermore, we provided an extensive ablation study to support our various design decisions.
I Introduction
@@
The manufacturing sector increasingly relies on robotics to address workforce shortages and enhance efficiency [1], particularly in repetitive and labor-intensive tasks. Traditionally, industrial robotics involves fixed robotic systems. A typical task is machine tending that automate the loading and unloading of materials and parts from production machines [2]. While effective, achieving this task with fixed setups is limited in flexibility and demands considerable human oversight.
@@
Advancements in robotics, driven by the imperatives of Industry 5.0 [3], suggest a pivotal shift from static to dynamic systems through the integration of mobile robotics. Such systems promise not only a higher degree of automation but also foster collaborative environments where robots and humans share and navigate the same workspace. For example, instead of a fixed robotic arm dedicated to a single machine, a fleet of mobile manipulators could autonomously navigate between multiple machines and storage areas, significantly boosting productivity and machine utilization.
@@
However, the deployment of multi-robot systems introduces complex challenges in coordination and control. Predominantly, solutions in the industry, such as those implemented in Amazon’s warehouses, have utilized a centralized approach where a single server orchestrates the tasks, trajectory planning, and collision avoidance of all robots, ensuring optimized fleet operations. This centralized model, while efficient, suffers from several disadvantages including single points of failure, high reliance on continuous communication, and limited adaptability in dynamic or unstructured environments [4].
@@
In contrast, decentralized robotic systems, where each unit operates independently and makes decisions in real-time, offer robust alternatives free from the constraints of centralized control. Reinforcement learning (RL), particularly with its recent advances in deep learning and computational capabilities supported by GPUs, presents a promising framework for such autonomous decision-making. RL enables robots to learn optimal behaviors through trial and error, using feedback from their environment to refine their actions [5]. Although extensively applied in areas like legged robot motion and fixed manipulator control, its application to mobile robot fleets, especially in complex tasks like machine tending, remains underexplored [6, 7].
@@
This paper introduces a novel reinforcement learning model tailored to the decentralized management of mobile robots for machine tending in manufacturing settings, aiming to leverage the full potential of autonomous systems in enhancing operational efficiency without the pitfalls of centralized control systems. In short, our main contributions are: 1)Propose and formulate a new more accurate and challenging scenario for multi-agent navigation for machine tending; 2) Design a novel dense reward function specifically for this scenario; 3) Design an attention-based encoding technique and incorporate it into a working baseline (AB-MAPPO); and 4) Provide an ablation study to verify the effect of various design decisions.
II Background and Related Work
II-A Machine Tending
@@
Despite the substantial research and development efforts in both academia and industry [8], the gap between automation demands and the available commercial products is still wide, and needs further work to be reduced.
@@
Due to the complexity of the machine tending problem, different researchers focused on different specific areas. For instance, Jia et al. [9] focused on recognizing the state of the machine by first detecting a specific computer numerical control (CNC) machine, and then using text detection and recognition techniques to get the status from its display.
@@
Al-Hussaini et al. [10] developed a semi-autonomous mobile manipulator that given a task from the operator generates the execution plan. It also estimates the risk of the plan execution and provides a simulation of the plan for the operator to approve it to be executed autonomously or to teleoperate the robot. In addition, Heimann et al. [2] presented a ROS2-based mobile manipulator solution with a focus on evaluating the system precision in reaching the machine location and reaching inside the machine. Moreover, Chen et al. [11] proposed a centralized method for multi-agent task assignment and path planing, where task assignment is informed by the task completion cost. Behbahani etl al. [12] used a learning-from-demonstration approach to teach a physical fixed robot to tend to one machine in a simplified tabletop setup.
@@
In a different direction, Burgess-Limerick et al. [13] developed a controller to allow a mobile manipulator to perform pick and place task on the move in a dynamic environment, reducing the overall task execution time.
II-B Reinforcement Learning
@@
RL has been explored for various robotic tasks such as pick and place, navigation, collision avoidance, locomotion and quadcopter control. [5, 14, 15]. Generally, RL methods can be divided into two main categories, on one hand, model-free techniques try to directly learn a policy that maximizes the returned reward. For instance, Proximal Policy Optimization (PPO) [16] is a very famous and widely used example, along with Deep Q-Networks (DQN) [17] and actor-critic methods [18]. On the other hand, model-based techniques try to explicitly learn a model for the environment dynamics and use that model to plan the actions directly [19], or to learn a policy by interaction with the learned dynamics model to reduce the needed interactions with the environment [20, 15].
@@
However, RL for machine tending is not a well-explored area of research, except for some works such as Iriondo et al. [7], where they used PPO and DDPG to teach a mobile manipulator to navigate to the location of a table to allow for the arm to pick objects from the table. Moreover, multi-agent reinforcement learning (MARL) is a growing research area that focuses on extending RL to multi-agent setups to expand its potential use. Two famous works in this area are MAPPO [21] and MADDPG [22] that extended PPO and Deep Deterministic Policy Gradient (DDPG) respectively.
@@
To the best of our knowledge, one of the only works presenting a MARL solution for machine tending is from Agrawal et al. [6]. This pioneering work introduces a multi-agent framework for integrated job scheduling and navigation for shop floor management exhibiting robustness under processing delays and failures. However, their implementation assumes the agents can accumulate parts from various machines without unloading them, as if they had infinite onboard storage. Distinct from prior models, our work addresses multi-agent, multi-machine tending by encompassing both the navigation to machines with ready parts and the transport of these parts to designated storage areas. This approach more accurately reflects real-world manufacturing scenarios, pushing MARL closer to full deployment on machine tending tasks. Moreover, our work can be integrated with low-level controller like  [13] to facilitate actual deployment in real robots.
III Problem Definition
@@
  • Navigation: Agents must navigate efficiently to and from machines, avoiding collisions with other agents and static obstacles while transporting parts to storage area.
@@
Object manipulation is acknowledged but not addressed in this study, focusing on the four constraints above. Key assumptions include instantaneous pick and place actions and external management of raw material supply, as supported by prior studies [13].
IV Methodology
IV-A MAPPO backbone
MAPPO, a prominent multi-agent reinforcement learning algorithm introduced by Yu et al. [21], operates on a dual-network architecture comprising an actor and a critic. The actor network, or policy, executes actions aimed at maximizing accumulated discounted returns, which are estimated by the critic based on the current state values.
@@ IV-A MAPPO backbone
IV-B Novel Attention-based Encoding for MAPPO
We introduce AB-MAPPO: an augment MAPPO’s architecture with the integration of a novel attention-based encoding mechanism into the critic network. This enhancement allows the actor to make decisions based on more accurately estimated values from the critic, thereby improving its performance indirectly as well. Figure 1 illustrates the design of our attention-based encoding mechanism for the critic. This design enables the critic to leverage all agents’ observations to develop a representation that encapsulates both spatial and temporal dynamics effectively.
@@ IV-B Novel Attention-based Encoding for MAPPO
The output from the multi-head attention module, based on the seminal paper of Vaswani et al. [23], is computed as follows:
@@
The attention module’s output is then flattened and concatenated with the original observations, and passed through a sequence of fully connected layers (FCs), normalization (Norm), a GRU, and a final FC layer to produce the value estimate. This layer configuration follows a standard, open-source implementation of MAPPO. For the actor, we retained the standard configuration, which includes a similar series of FCs, Norms, and a GRU layer, differing only in the final output layer, which matches the number of possible actions.
IV-C Observation Design
In addressing the multi-agent, multi-machine tending problem, it was essential to design an efficient observation scheme that captures essential information while excluding superfluous details. After extensive empirical testing to determine the optimal data inclusion by trying different observation options (more about it in section VI), each agent’s observation was crafted to include:
@@ 4. Time Penalty ($R_{t}$):
V Experiments
V-A Simulation Setup
For our experiments, we leveraged the Vectorized Multi-Agent Simulator (VMAS) [24], specifically designed for efficient multi-agent reinforcement learning (MARL) benchmarking. VMAS integrates a 2D physics engine and supports the well-known Gym simulation environments interface. Its architecture allows for straightforward customization and expansion, facilitating the development of new and more complex scenarios. More about the action space and the simulator in the supplementary materials.
Figure 2 illustrates the multi-agent multi-machine tending scenario we developed. In this setup, three agents, depicted in red, initiate from their designated idle positions at the top of the environment. The production machines are highlighted in green, with machine blockers in gray obstructing direct access to the machines from one side. The storage area is marked in blue, and the walls outlining the environment are shown in black.
Figure 2: Multi-agent (red) multi-machine (green) tending scenario designed in VMAS, including obstacles (gray) and storage area (blue).
@@ 3. Resource Utilization Factors:
V-C Results and Adaptability
Table I presents a comparative analysis of AB-MAPPO against the standard MAPPO. The data clearly demonstrate that AB-MAPPO significantly surpasses MAPPO across all evaluated metrics. Notably, it achieves a 40% reduction in collisions, enhancing safety. Furthermore, it exhibits a 16% increase in parts collection and a 20% improvement in parts delivery. Additionally, the model boosts average machine utilization and agent utilization by 8% each, indicating a more efficient use of resources.
TABLE I: Evaluation results for our model compared to MAPPO: Mean (stddev).
@@ V-C Results and Adaptability
However, significantly increasing the distance between the machines and the storage area adversely affects performance. We attribute this decline to the increased challenge agents face in locating the storage area after part collection, necessitating exploration over a larger area. The placement of the storage area—whether centrally or peripherally—does not markedly influence performance. Nevertheless, starting agents in close proximity to one another can impede their learning due to an increased incidence of collisions at the outset, which the collision penalty exacerbates by discouraging exploration. Additional experiments and results are provided in the supplementary materials [^1] https://initrobots.ca/mammt/.
VI Ablation Study
VI-A Experiments on Observation Design
To develop an effective observation strategy for our model, we initially assessed various observation configurations using MAPPO to determine their impact on performance. Table II presents a baseline experiment (Exp1), which incorporates all observation components detailed in IV-C including all agents’ velocities and the relative positions of the walls.
@@ VI-A Experiments on Observation Design
TABLE II: Evaluation results for various observation components: Mean (stddev).
@@
Furthermore, enhancements in parts delivery correlated well with improvements in both machine and agent utilization rates. This suggests that factors contributing positively to operational efficiency also tend to optimize resource usage. performance. Utilizing the same base experiment setup as before, all reward components outlined in (15) were initially included, along with reward sharing where all agents receive the pick or place reward if at least one agent successfully picks up or places a part, respectively. Subsequent experiments each isolate and modify a specific reward feature to analyze its effects. Where Exp2 eliminates the time penalty, Exp3 removes reward sharing, Exp4 omits the uncollected parts penalty, Exp5 applies an increasing uncollected parts penalty as described in the latter part of (13), and Exp6 removes the distance-based reward. Table III details these comparisons.
TABLE III: Evaluation results for various reward components: Mean (stddev).
@@
Additionally, we assessed the performance of combining the most effective observation and reward options. Table IV presents these results. Not all combinations led to performance enhancements. For instance, Exp1, which combines removing velocity observation (RV) with observing machine blockers’ positions (OMB), did not improve outcomes. However, Exp2, which combines RV with no reward sharing (NSH), showed enhanced performance, as did Exp3, which pairs no velocity observation (NV) with two-corner entity representation (COR). Contrarily, combining all selected options (NV, COR, OMB, and NSH) in Exp4 did not yield better results. Ultimately, the setup from Exp2 was chosen as the optimal configuration for our observation and reward strategy.
TABLE IV: Combining the best performing options: Mean (stddev).
@@
VII Conclusion
This study significantly advances multi-agent multi-machine tending by addressing unexplored complexities. We enhanced the MAPPO architecture with AB-MAPPO, incorporating a novel attention-based encoding to improve feature representation significantly. These innovations establish a new baseline and pave the way for future advancements. Rigorously tested across diverse environments and supported by a comprehensive ablation study, AB-MAPPO demonstrated its efficacy and adaptability. Looking forward, we plan to extend these innovations to more dynamic scenarios, including real-time material feeding and interactive object manipulation, to bridge the gap between theoretical research and practical applications.

### v2

@@ Abstract
Robotics can help address the growing worker shortage challenge of the manufacturing industry. As such, machine tending is a task collaborative robots can tackle that can also highly boost productivity. Nevertheless, existing robotics systems deployed in that sector rely on a fixed single-arm setup, whereas mobile robots can provide more flexibility and scalability. We introduce a multi-agent multi-machine tending learning framework by mobile robots based on Multi-agent Reinforcement Learning (MARL) techniques with the design of a suitable observation and reward. Moreover, we integrate an attention-based encoding mechanism into the Multi-agent Proximal Policy Optimization (MAPPO) algorithm to boost its performance for machine tending scenarios. Our model (AB-MAPPO) outperforms MAPPO in this new challenging scenario in terms of task success, safety, and resource utilization. Furthermore, we provided an extensive ablation study to support our design decisions.
Introduction
@@
The manufacturing sector increasingly relies on robotics to address workforce shortages and enhance efficiency (Kugler 2022), particularly in repetitive and labor-intensive tasks. Traditionally, industrial robotics involves fixed robotic systems. A typical task is machine tending that automates the loading and unloading of materials and parts from production machines (Heimann et al. 2023). While effective, achieving this task with fixed setups is limited in flexibility and demands considerable human oversight.
@@
Instead of a fixed robotic arm dedicated to a single machine, a fleet of mobile manipulators could autonomously navigate between multiple machines and storage areas, significantly boosting productivity and machine utilization.
@@
However, the deployment of multi-robot systems introduces complex challenges in coordination and control. Predominantly, solutions in the industry, such as those implemented in Amazon’s warehouses, have used a centralized approach where a single server orchestrates the tasks, trajectory planning, and collision avoidance of all robots, ensuring optimized fleet operations. This centralized model, while efficient, suffers from several disadvantages including single points of failure, high reliance on continuous communication, and limited adaptability in dynamic or unstructured environments (Fan et al. 2018).
@@
In contrast, decentralized robotic systems, where each unit operates independently and makes decisions in real-time, offer robust alternatives free from the constraints of centralized control. Reinforcement learning (RL), particularly with its recent advances in deep learning and computational capabilities supported by GPUs, presents a promising framework for such autonomous decision-making. RL enables robots to learn optimal behaviors through trial and error, using feedback from their environment to refine their actions (Singh, Kumar, and Singh 2022).
Despite being a very promising technique, most of the works that investigate RL for multi-agent tasks, do it for games like Starcraft Multi-Agent Challenge (SMAC) (Guo et al. 2024; Zhou et al. 2021) and Soccer (Lobos-Tsunekawa, Srinivasan, and Spranger 2022) or simplified scenarios such as goal coverage and formations (Agarwal, Kumar, and Sycara 2019; Long et al. 2020; Chen et al. 2020) distancing MARL from deployable real-world scenarios.
@@
In contrast, this paper introduces a novel reinforcement learning model tailored to a realistic scenario of decentralized management of mobile robots for machine tending in manufacturing settings autonomously handling both task assignment and navigation. Aiming to leverage the full potential of autonomous systems in enhancing operational efficiency. In short, our main contributions are: 1)Propose and formulate a new more realistic, and challenging scenario for multi-agent navigation for machine tending; 2) Design a novel dense reward function specifically for this scenario; 3) Design an attention-based encoding technique and incorporate it into a working baseline (AB-MAPPO). We also include an ablation study to verify the effect of various design decisions.
Background and Related Work
Machine Tending
@@
Despite the substantial research and development efforts in both academia and industry (Li et al. 2023), the gap between automation demands and the available commercial products is still wide and needs further work to be reduced.
@@
Due to the complexity of the machine tending problem, different researchers focused on different specific areas. For instance, Jia et al. (Jia et al. 2022) focused on recognizing the state of the machine by first detecting a specific computer numerical control (CNC) machine and then using text detection and recognition techniques to get the status from its display.
@@
Al-Hussaini et al. (Al-Hussaini et al. 2020) developed a semi-autonomous mobile manipulator that given a task from the operator generates the execution plan. It also estimates the risk of the plan execution and provides a simulation of the plan for the operator to approve it to be executed autonomously or to teleoperate the robot. In addition, Heimann et al. (Heimann et al. 2023) presented a ROS2-based mobile manipulator solution with a focus on evaluating the system precision in reaching the machine location and reaching inside the machine. Moreover, Chen et al. (Chen et al. 2021) proposed a centralized method for multi-agent task assignment and path planning, where task assignment is informed by the task completion cost. Behbahani et al. (Behbahani et al. 2021) used a learning-from-demonstration approach to teach a physical fixed robot to tend to one machine in a simplified tabletop setup.
@@
In a different direction, Burgess-Limerick et al. (Burgess-Limerick et al. 2024) developed a controller to allow a mobile manipulator to perform pick and place task on the move in a dynamic environment, reducing the overall task execution time.
Reinforcement Learning
@@
RL has been explored for various robotic tasks such as pick and place, navigation, collision avoidance, locomotion, and quadcopter control. (Singh, Kumar, and Singh 2022; Zhu and Zhang 2021; Wu et al. 2023). Generally, RL methods can be divided into two main categories, on one hand, model-free techniques try to directly learn a policy that maximizes the returned reward. For instance, Proximal Policy Optimization (PPO) (Schulman et al. 2017) is a very famous and widely used example, along with Deep Q-Networks (DQN) (Mnih et al. 2013) and actor-critic methods (Konda and Tsitsiklis 1999). On the other hand, model-based techniques try to explicitly learn a model for the environment dynamics and use that model to plan the actions directly (Hafner et al. 2019), or to learn a policy by interaction with the learned dynamics model to reduce the needed interactions with the environment (Sutton 1991; Wu et al. 2023).
However, RL for machine tending is not a well-explored area of research, except for some works such as  (Iriondo et al. 2019), where they used PPO and DDPG to teach a single mobile manipulator to navigate to the location of a table that has the object to pick. Despite targeting one agent and one machine setup, their work is still far from being deployable in reality because it does not address part dropping, learning to navigate when the part is ready, or navigating multiple times to pick and place multiple parts.
@@
On the other hand, multi-agent reinforcement learning (MARL) is a growing research area that focuses on extending RL to multi-agent setups to expand its potential use. Two famous works in this area are MAPPO (Yu et al. 2021) and MADDPG (Lowe et al. 2017) that extended PPO and Deep Deterministic Policy Gradient (DDPG) respectively.
@@
To the best of our knowledge, one of the only works presenting a MARL solution for machine tending is (Agrawal et al. 2021). This pioneering work introduces a multi-agent framework for integrated job scheduling and navigation for shop floor management exhibiting robustness under processing delays and failures. However, their implementation assumes the agents can accumulate parts from various machines without unloading them, as if they had infinite onboard storage, making the solution non-deployable in the real world.
Distinct from prior models, our work addresses multi-agent, multi-machine tending by encompassing task assignment, the navigation to machines with ready parts, and the transport of these parts to designated storage areas. This approach more realistically reflects real-world manufacturing scenarios, pushing MARL closer to full deployment on machine tending tasks. Moreover, our work can be integrated with low-level controllers like  (Burgess-Limerick et al. 2024) to facilitate actual deployment in real robots.
Problem Definition
@@
  • Navigation: Agents must navigate efficiently to and from machines, avoiding collisions with other agents and static obstacles while transporting parts to the storage area.
@@
Object manipulation is acknowledged but not addressed in this study, focusing on the four constraints above. Key assumptions include instantaneous pick and place actions and external management of raw material supply, as supported by prior studies (Burgess-Limerick et al. 2024).
Methodology
MAPPO backbone
MAPPO, a prominent multi-agent reinforcement learning algorithm introduced by Yu et al. (Yu et al. 2021), operates on a dual-network architecture comprising an actor and a critic. The actor network, or policy, executes actions aimed at maximizing accumulated discounted returns, which are estimated by the critic based on the current state values.
@@
Novel Attention-based Encoding for MAPPO
We introduce AB-MAPPO: an upgraded MAPPO’s architecture with the integration of a novel attention-based encoding mechanism into both actor and critic networks. This enhancement allows the actor to make decisions based on more accurately estimated values from the critic, thereby improving its performance indirectly as well. Figure 1 illustrates the design of our attention-based encoding mechanism for the critic. This design enables the critic to leverage all agents’ observations to develop a representation that encapsulates both spatial and temporal dynamics effectively.
@@
The output from the multi-head attention module, based on the seminal paper of Vaswani et al. (Vaswani et al. 2017), is computed as follows:
@@
The attention module’s output is then flattened and concatenated with the original observations and passed through a sequence of fully connected layers (FCs), normalization (Norm), a GRU, and a final FC layer to produce the value estimate. This layer configuration follows a standard, open-source implementation of MAPPO. For the actor, we retained the standard configuration, which includes a similar series of FCs, Norms, and a GRU layer, differing only in the final output layer, which matches the number of possible actions.
Observation Design
In addressing the multi-agent, multi-machine tending problem, designing an efficient observation scheme that captures essential information while excluding superfluous details was crucial. After extensive empirical testing to determine the optimal data inclusion by trying different observation options (more about it in section Ablation Study), each agent’s observation was crafted to include:
@@ 4. Time Penalty ($R_{t}$):
Experiments
Simulation Setup
For our experiments, we leveraged the Vectorized Multi-Agent Simulator (VMAS) (Bettini et al. 2022), specifically designed for efficient multi-agent reinforcement learning (MARL) benchmarking. VMAS integrates a 2D physics engine and supports the well-known Gym simulation environments interface. Its architecture allows for straightforward customization and expansion, facilitating the development of new and more complex scenarios.
Figure 2 illustrates the multi-agent multi-machine tending scenario we developed. In this setup, three agents, depicted in red, initiate from their designated idle positions at the top of the environment. The production machines are highlighted in green, with machine blockers in gray obstructing direct access to the machines from one side. The storage area is marked in blue, and the walls outlining the environment are shown in black. While dotted lines in the figure point to the actual robot(our RanGen robot composed of a Kinova Gen3 arm on top of an AgileX Ranger Mini mobile base), machines (CNC Universal Milling Machine DMU 50) and storage shelves that we are planning to use for real deployment.
Figure 2: Multi-agent (red) multi-machine (green) tending scenario designed in VMAS, including obstacles (gray) and storage area (blue), with dotted lines pointing to the actual robot (our RanGen robot composed of a Kinova Gen3 arm on top of an AgileX Ranger Mini mobile base), machines (CNC Universal Milling Machine DMU 50) and storage shelves that we are planning to use for real deployment.
In VMAS agents’ actions are represented as physical forces $f_{x}$ and $f_{y}$. The physics engine takes into account this input force and other factors like the velocity damping, gravity, and forces due to the collisions with other objects (Bettini et al. 2022). We used 5 discrete actions 0: no force (translated internally to $f_{x}=0,f_{y}=0$), 1: Leftward acceleration ($f_{x}=-1,f_{y}=0$), 2: rightward acceleration ($f_{x}=1,f_{y}=0$), 3: downward acceleration ($f_{x}=0,f_{y}=-1$) and 4: upward acceleration ($f_{x}=0,f_{y}=1$). This holonomic motion is also supported by our intended mobile base (the AgileX Ranger Mini).
@@ 3. Resource Utilization Factors:
Results and Adaptability
Table 1 presents a comparative analysis of AB-MAPPO against the standard MAPPO. The data clearly demonstrate that AB-MAPPO significantly surpasses MAPPO across all evaluated metrics. Notably, it achieves a 40% reduction in collisions, enhancing safety. Furthermore, it exhibits a 16% increase in parts collection and a 20% improvement in parts delivery. Additionally, the model boosts average machine utilization and agent utilization by 8% each, indicating a more efficient use of resources.
@@
Table 1: Evaluation results for our model (AB-MAPPO) compared to MAPPO: showing the Mean and (standard deviation) of the number of collected parts, delivered parts, collisions, and average machine and agent utilization.
@@
However, significantly increasing the distance between the machines and the storage area adversely affects performance. We attribute this decline to the increased challenge agents face in locating the storage area after part collection, necessitating exploration over a larger area. The placement of the storage area—whether centrally or peripherally—does not markedly influence performance. Nevertheless, starting agents in close proximity to one another can impede their learning due to an increased incidence of collisions at the outset, which the collision penalty exacerbates by discouraging exploration. Additional experiments and results are provided in the project website.
Ablation Study
Experiments on Observation Design
To develop an effective observation strategy for our model, we initially assessed various observation configurations using MAPPO to determine their impact on performance. Table 2 presents a baseline experiment (Exp1), which incorporates all observation components detailed in Observation Design including all agents’ velocities and the relative positions of the walls.
@@
Table 2: Evaluation results for various observation components: showing the Mean and (standard deviation) of the number of collected parts, delivered parts, collisions, and average machine and agent utilization.
@@
Furthermore, enhancements in parts delivery correlated well with improvements in both machine and agent utilization rates. This suggests that factors contributing positively to operational efficiency also tend to optimize resource usage.
Experiments on Reward Design
This section continues our evaluation by focusing on the different components of the reward structure to determine their individual impacts on performance. Utilizing the same base experiment setup as before, all reward components outlined in (15) were initially included, along with reward sharing where all agents receive the pick or place reward if at least one agent successfully picks up or places a part, respectively. Subsequent experiments each isolate and modify a specific reward feature to analyze its effects. Where Exp2 eliminates the time penalty, Exp3 removes reward sharing, Exp4 omits the uncollected parts penalty, Exp5 applies an increasing uncollected parts penalty as described in the latter part of (13), and Exp6 removes the distance-based reward. Table 3 details these comparisons.
@@
Table 3: Evaluation results for various reward components: showing the Mean and (standard deviation) of the number of collected parts, delivered parts, collisions, and average machine and agent utilization.
@@
Additionally, we assessed the performance of combining the most effective observation and reward options. Table 4 presents these results. Not all combinations led to performance enhancements. For instance, Exp1, which combines removing velocity observation (RV) with observing machine blockers’ positions (OMB), did not improve outcomes. However, Exp2, which combines RV with no reward sharing (NSH), showed enhanced performance, as did Exp3, which pairs no velocity observation (NV) with two-corner entity representation (COR). Contrarily, combining all selected options (NV, COR, OMB, and NSH) in Exp4 did not yield better results. Ultimately, the setup from Exp2 was chosen as the optimal configuration for our observation and reward strategy.
@@
Table 4: Combining the best performing options: Mean (stddev).
Conclusion
This study applies MARL to a challenging real-world industrial scenario (machine tending) aiming to leverage the full potential of multiple mobile robots. We enhanced the MAPPO architecture with AB-MAPPO, incorporating a novel attention-based encoding to improve feature representation significantly. These innovations establish a new baseline and pave the way for future advancements. Rigorously tested across diverse environments and supported by a comprehensive ablation study, AB-MAPPO demonstrated its efficacy and adaptability. Looking forward, we plan to extend these innovations to more dynamic scenarios, including real-time material feeding and interactive object manipulation, to bridge the gap between theoretical research and practical applications.